Scalability analysis of low-level image processing

algorithms

implemented on a 2-dimensional array of transputers

Zoltan Juhasz

University of Veszprem, Hungary

Abstract

Transputers have been used for many years with varying degree of success to build parallel image processing systems. While many experimental and prototype systems have been developed, the effect of the transputer systems’ modularity, flexibility and extendibility on performance has been overlooked.

This talk focuses on how 2-dimensional transputer arrays can be used to implement scalable, parallel image processing algorithms. The different factors affecting overhead are analyzed and then data transmission and communication are investigated. Various communication methods are compared with respect to scalability and the most suitable method is used to implement image processing algorithms. Once the overhead time can be characterized, the scalability of some representative operations are analyzed to predict the performance and speedup of a ”P” processor system.

The goal in examining these critical points is to provide guidelines for system and algorithm design. This will help designers create scalable image processing systems.